North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6



14th to 20th January 2021

On Tuesday 19th, it was raining all day, but a lot milder, for the Q6 working party of Bill Dobson, Ian Pearson and Jon Bradley, who had plenty to do in Deviation Shed. Bill was busy gapping the piston rings - a long and laborious job, cutting and filing to achieve the 1/8” gap required. Ian and Jon spent the day adjusting the pipe work so that the unions connected squarely, and fitting a bracket onto the steam heat and vacuum pipes.

It was still raining on Wednesday, with Ed Bolam joining Bill Dobson and Ian Pearson. Bill first went down to the MPD to check on progress with the Q6 cross heads: after re-metalling, a start had been made on the necessary machining. Bill then spent the rest of the day completing the gapping of the piston rings. After Ed had cleaned and lit the stove, he set up a bench with Ian at the south end of Deviation shed, and dressed and primed the left hand cab shoulder cladding sheet. However, he only dressed the one piece due to the amount of dust produced while removing the old paint. Definitely an outside job for a fine day! He also cleaned and oiled the slide bars, and oiled inside the left hand cylinder. Meanwhile, Ian fitted the right hand side rod securing nuts and taper pins.

 

7th to 13th January 2021

A full week of activity to report -

7 January: Q6 - Ian Pearson and Bill Dobson met at lunchtime at Deviation Shed with the boiler in steam down at the MPD. However, Chris Kelly informed them just after lunch that he had had to fail the in house steam test due to the tea tray studs leaking. Once the boiler had cooled they would be removed. The insurance out of frames steam test arranged for the Monday was therefore cancelled and the boiler inspector informed.

12 January: Q6 - it was a cold frosty morning at Grosmont on Tuesday for Bill Dobson, Nigel Bill, Ian Pearson and Sean Bowler. Jon Bradley joined later in the afternoon. Gordon Wells also called in to rewire the kettle and water heater onto separate circuit breakers so they don’t trip out.

Bill removed the steam supply valve that feeds the vacuum brake. This was leaking steam and he found the seat required cutting. This job took most of the day, because, as is often the case, searching for cutters was a nightmare. In the end, one was borrowed from the MPD. A new brass washer is being made in place of a worn out one, so the valve will be re-fitted shortly. The water had been drained out of the boiler on Saturday morning. Nigel and Ian removed the pressure gauge and gauge glass protectors. In the afternoon, Sean completed the slide bar alignment then took both crossheads down to the MPD for measuring and checking. These will be re-metalled and machined. Later in the afternoon, with Jon assisting, the vacuum pipe was fitted.

An inspection of NELPG ladders was carried out, all of which are in Deviation Shed. One long, old, ladder has been scrapped due to defective rungs, and another has been labelled 'Not to be Used' as the integrated stabiliser has been bent and needs to be repaired. The footpath between Nos 8 and 7 road had become very cluttered with Items of the Q6 making the emergency door difficult to access. The items have been removed and the footpath is now clear.

Down at the MPD, tons of silt and ash was removed from drain sumps at the north end outside the MPD by a contract firm. We have been told that the water from Deviation Shed pit will have to be diverted away from the ash pit outside our building directly to the main drain to stop the ash getting into the traps and dumps.

13 January: Q6 - was another very cold but fine day with Bill Dobson, Ed Bolam, Ian Pearson and a late morning arrival of Nigel Bill who had picked up some consumables from Arco - 3 boxes of nitrile disposable gloves and 1 box of disposable face masks.

After getting the stove fired up, Bill, Ed and Ian fitted the steam heat pipes. They still require some securing, but are in position and also out of the way of the footpath. When Nigel arrived he spent most of the day fitting split pins to the slide bar bolts on both sides. Bill, Ed and Ian fitted fitted the front control sands rod and the left to right sands operating rod which connects to the front control rod. Ed also cleaned the the gauge glass protectors and replaced a spring on one of the protectors. Then he fitted the cylinder cocks, although some more fine adjustments are required.

There was also some tidying up in the workshop. An empty red cabinet, which has been in the way upstairs, was brought down and now sits on the bench at the south end of the workshop with a variety of tools stored in it.

Down at the MPD, a train was marshalled comprising of a well wagon with the boiler of 44806 secured to it and a brake van and Class 25 ready to move to Newbridge yard on Thursday. The boiler is going Riley engineering.

The next Q6 working parties will be on Tuesday 19 January with Ian Pearson and Bill Dobson, and Wednesday 20 January with Ian Pearson, Bill Dobson and Nigel Bill.

 

4th to 6th January 2021

This first report of the New Year covers three days. After Bill and Ian had discussed the work required to prepare for the insurance steam test on Monday 11 January, they decided to work at Grosmont on Monday this week, as well as Tuesday and Wednesday.

So, on a cold Monday morning, with heavy wintery showers, Bill and Ian met Sean Bowler at Deviation Shed to discuss what needed to be done with the tight gradient pin and bush on the leading section left hand side rod. After removing the pin which Sean ground slightly, he then dressed the inside of the bush with a small abrasive wheel. Bill and Ian started work setting up jacks under the foundation ring in preparation for lifting, to allow removal of the wooden beam and its replacement with the steel beam Bill acquired recently. But the steel beam had first to be reinforced by welding steel plate into the web of the beam, so wouldn't be ready for another 24 hours. By lunchtime Sean had finished dressing the bush. The gradient pin was tried inside the bush and was a good fit. Afterwards the front section of side rod was assembled, and the side rods fitted back onto the wheels.

Tuesday was another cold wintry day with Bill Dobson, Nigel Bill and Ian Pearson rostered, along with Sean Bowler and Dave Brown. After discussion with Sean and Dave, the main job for Sean was to prepare for lining up the slide bars, but this would mean the Q6 frames could not be moved. So it was decided that Sean would drive the tele handler and take the pallet of bagged shot blast material out of Deviation Shed to covered accommodation in a metal container near the oil store. After this, the Q6 was shunted outside and the steam reverser was taken by the tele handler to the south end of Deviation Shed and fitted back onto the Q6 frames which were then shunted back into Deviation Shed. Sean and Dave were then able to start setting up their equipment for lining up the slide bars. Unfortunately, Sean O'Driscoll's law continues to apply in 2021, and some of the fittings were found to be not compatible with the Q6: Sean had to make some new fittings. While that was being done, Dave fitted some split pins on the horn guide bolts that never got done when the frames were down at the MPD. Bill and Ian started securing the steam reverser which they planned to complete the next day. After lunch, Bill, Ian and Nigel set about removing the wooden beam from under the firebox foundation ring (see photo) using strops and a pull lift. Eventually it was replaced with a steel beam (see photo).

63395 wooden beam being removed from under the foundation ring by Nigel Bill and Bill Dobson on 5 January 2021 - Ian Pearson

63395 replacement steel beam fitted under the foundation ring on 5 January 2021 - Ian Pearson

It was no warmer on Wednesday for Bill Dobson, Ian Pearson and Ed Bolam, again, along with Sean Bowler and Dave Brown. Sean and Dave spent the day working on the left hand slide bar alignment (see photo). Bill also discussed the alignment with Sean and established that the cross heads will also have to be re-metalled. Bill has measured the boiler barrel and firebox area for insulation which will be ordered shortly, and also for new flexible axle box oil pipes, of which 16 will be required, 18” long. Ian and Ed completed fitting the steam reverser securing bolts and split pinning. In the morning, Chris Kelly fitted the bottom left hand mud hole door. This had been removed to let the boiler water out on the Tuesday before Christmas. The boiler was then filled half full and after lunch Ed and Ian set, and lit, a warming fire in the Q6. Ed also fitted both cab bottom steps which were removed when the frames were on No 4 Road in the MPD to allow the jacks to be fitted when the frames were lifted.

63395 with Sean Bowler and Dave Brown lining up the left hand slide bars on 6 January 2021 - Ian Pearson

 

22nd December 2020

Ian Pearson reports that it was a cold and frosty start for the last working party of 2020 at Grosmont yesterday, comprising Bill Dobson, Nigel Hall, Nigel Bill and Ian himself.

Nigel H got the stove going, which was very welcoming on this cold morning, before he started painting. All four eccentric rods have now had a coat of red gloss, and he also did some touching up between the frames. Nigel B completed the monthly and yearly defibrillator checks.

The main job of the day though was to move the Q6 frames from the rear of No 4 road at the MPD back into Deviation Shed. Ian spent most of the morning shunting with Charlie Wood and Chris Cubitt driving the 04 shunter D2207. By lunchtime the Q6 frames were moved into No 7 Road in Deviation Shed, with the B1 in front and 5428 Eric Treacy moved to the wheel drop shed.

Nigel B and Ian then cleaned all the old cardboard from the 4 foot area where the Q6 frames stood, which had been used on the oily floor while everyone worked between the frames, to keep them a little cleaner.

Gordon Wells, who was working down at the MPD, came in and repaired a couple of lights in Deviation Shed above the West side roller shutter door. 2 fluorescent lights were repaired, one with a new tube and the other with a new starter. Two roof lights near the stove area have been out for a while, and Gordon found that they had only tripped out. The circuit breaker was switched back in and they are now all working OK.

63395 with left hand rods removed in Deviation Shed on 22 December 2020 - Ian Pearson

After lunch, a concentrated effort by all four saw the removal of the left hand side rods (see photo) to allow the removal of the leading gradient pin which was found to be tight in the bush. The pin was eventually removed, but the bush, which was proud and also required removal, proved very stubborn, and, after trying various methods of extraction, by late afternoon it was decided to leave it until 5 January 2021 when the next Working Party is scheduled, involving Bill Dobson, Ian Pearson, Nigel Bill, and Nigel Hall, followed by Wednesday 6 January with Bill Dobson, Ian Pearson, Steve Hyman and and Ed Bolam, coronavirus and tier restrictions permitting.

In the meantime, Bill has arranged with the boiler inspector for the out of frames steam test to be held on 11 January at 1pm.

Following the retirement of Bryan Orange as Deviation Shed caretaker after 10 years in charge (he is taking on the Chairmanship of the HRA member services Training and Development Committee in the New Year), I can report that Ian Pearson and Nigel Bill have volunteered to take up the mantle of responsibility as joint caretakers for the building. My thanks to them for so doing, and also to Bryan Orange for all the hard work he has put in over the years in looking after the building and making numerous improvements to its fabric and facilities from which we are all benefitting. He will be continuing with the NELPG JVs, but I wish him good luck with his new role.

At the MPD, Repton was on a steam test and repairs to the J27's valve covers were being carried out. The No 5 Road roller shutter door was also being repaired by N.E.S.S. following a losing argument with a locomotive. Tomorrow sees the last of the Christmas Re-imagined services from Pickering, and the four Moorlander dining trains scheduled for 31 December, and 1 to 3 January are also due to run, again, subject to any further changes to the coronovirus controls which seem to be occurring on almost a daily basis. However, the planned Optimist and Rail Trail services planned between Boxing Day and 3 January have been cancelled, and whether services operate after that date will be subject to review and decision every two weeks. Regardless, the NYMR is hoping to run trains over the February half term (13 - 21 February) with a possible special visitor. So the message is to keep an eye on the NYMR website for up to date details of what is, and is not, running in early 2021.

Finally, it just remains for me to wish you all, and in particular my special correspondent Ian Pearson, all the very best for Christmas and the New Year. 2020 has been the most unimaginable and difficult one I have ever known, so my thanks to all of you, including our NYMR colleagues in the MPD in particular, for your contribution and support, particularly over the last nine months. Although 2021 looks as if it is going to start in similar vein, let's hope that the arrival of the vaccines will herald the beginning of the end of the pandemic and a return, eventually, to a more normal life. As ever though, until that time comes, please do continue to take care and, most importantly, stay safe.

 

15th and 16th December 2020

This report really starts last Thursday (10 December), which saw the Q6 re-wheeled and all eight horn stays fitted by Sean Bowler and Dave Brown, with assistance from Nigel Bill who was on a cleaning turn for the NYMR and Ian Pearson. Since then, on Friday and Monday, Sean fitted all the springs.

On a fine mild Tuesday, Bill Dobson, Nigel Bill, Jon Bradley and Ian Pearson then made a concerted effort in fitting all the brake hangers and tables, plus all the brake rods.

Wednesday though was a miserable wet mild day for Steve Hyman, Ed Bolam, Nigel Hall, Bill Dobson and Ian Pearson. Nigel Hall spent the day painting, putting another coat of black gloss on the backplate and the steam heat pipes, whilst after lunch he rubbed down the four eccentric rods and gave them a coat of red oxide.

The rest of the team made another collective effort with Sean and Dave, splitting each set of side rods into their three sections up at the wheel drop shed. They were then dragged outside so Sean could pick them up with the tele handler and take them to the MPD where they were off loaded onto a trolley and moved to the rear of No 4 Road for reassembly and fitting. The left side was fitted first, then, after lunch, the righthand side was fitted with extra help from Chris Cubitt. This was no easy job with lots of lifting, pulling, dragging, and lifting various pieces of timber packing and moving the jacks, along with pinch barring the locomotive to line up the wheels. Eventually, by late afternoon the objective for the day was achieved and the fitting of the side rods completed. In one final task for the day, the water was drained from the Q6 boiler as the boiler inspector visit for the out of frames steam test is now provisionally arranged for 11 January 2021.

On Tuesday and Wednesday, Lucie was on steam heat duty and the S15 on the Moorlander dining train. Meanwhile, following Repton failing the previous week, a warming fire was put in the J27 on Tuesday ready for a successful steam test on Wednesday and it was in charge of the Moorlander dining train today in glorious sunshine and will be again tomorrow - but in sunshine again?

 

8th and 9th December 2020

On a cold, misty, drab, Tuesday morning at Grosmont, Bill Dobson, Nigel Hall, Nigel Bill and Ian Pearson arrived to find smoke issuing from the Q6 chimney. A warming fire was lit on Monday afternoon then relit on Tuesday morning (see photo).

63395 boiler out of frames steam test on 8 December 2020 -  Nigel Hall

Away from that excitement, Nigel B and Ian continued to clean between the frames. Nigel B also dressed some flaking paint from the weigh shaft area and applied a coat of red oxide on the bare metal. The lifting jacks had arrived back from Pickering on Monday, but, when tested, they would not work. An electrical fault was found in the control unit which Chris White (electrician) rectified during the morning. So, before lunch, all four jacks were moved into position by Nigel B and Ian, Chris Cubitt, Sean Bowler, Andrew Jeffery and Dave Cholmondeley. Moving these jacks is very heavy work, but they were now set ready for Sean to do some final adjustments and measurements to the horn guides the next day. Bill started rubbing down the vacuum pipes which Ed cleaned last week, while Nigel H gave the back plate, cab steps and tea tray a coat of black gloss paint. Later Nigel H dressed the steam heat pies with a flapper wheel and gave them a coat of black primer as well as giving the vacuum pipes a coat of red oxide where required.

After lunch the steam test on the boiler had to be abandoned for various reasons, and which were to be sorted out the next day. These included the absence of blanks on the header: these were found upstairs in the Deviation Shed workshop. Ian made some new gaskets for them and they were to be fitted tomorrow. Also the packing under the foundation ring area required some alteration, and the blower pipe from the manifold to the operating valve required fitting. Bill started this job just before home time and was to continue the next day. It now looks as if the boiler inspector will be unable to come for the formal out of frames steam test until January 2021.

The weather on Wednesday was more of the same, although it brightened up later for Steve Hyman, Nigel Hall, Bill Dobson, Ian Pearson and, at lunchtime, Jonathan Bradley.

Steve fitted the blower supply pipe from the manifold to the blower valve on the back head and Ian fitted the two blanks to the header in the smokebox. One stud was missing so Bill asked Barry Nesom to make a new one. This was fitted and both blanks are now in situ on the header. Ian removed the gauge frame protectors and these are now stored in a black plastic bucket in the large locker upstairs in Deviation shed. They require some attention - new springs,cleaning and painting. Nigel H continued to paint - another coat of black gloss on the steam heat pipes, red gloss on the vacuum pipes, and eventually got down to the frames later in the afternoon. He then painted all the parts that had been painted in red oxide yesterday with red gloss.

After lunch Steve and Jon worked outside on back head parts, dressing by removing old flaking paint but there is still quite an amount to do so Nigel H is going to acquire some dressing disks.

Bill went to local steel stockists to buy a steel beam of 6 feet in length for fitting underneath the foundation ring as the wooden beam started to burn yesterday. Bill is arranging delivery to Grosmont sometime shortly. The water in the boiler is to be drained on Friday, then it will be jacked up, and the wooden beam replaced with the steel beam. Sean and Dave fitted the intermediate and leading wheel sets with new under keep pads. The four standing jacks were removed. Some final measurement checks on the horn guides showed all to be OK. The wheels are to be fitted back into the frames today (Thursday).

Locomotives in steam on both days were Lucie on steam heat duties and S15 on the Moorlander.

 

1st and 2nd December 2020

Well here we are at December 1st and Christmas and the New Year are just around the corner. But on a fine cold Tuesday morning at Grosmont, Nigel Hall, Bill Dobson, Nigel Bill and Ian Pearson arrived for the first of the week's Q6 working parties.

Last week Terry was repairing copper pipes. However, one was forgotten - the steam pipe from the steam brake to the brake cylinder which has the oil supply pot attached. It was noted when removed that the pipe was cracked about 20 inches below the oil pot. But on initial inspection in the morning, Ian and Nigel B could not find the crack, so the pipe was returned back upstairs. Sean O'Driscoll's law applies at Grosmont though, and, later in the afternoon, Ian found the crack on the steam brake pipe while looking for another item in the upstairs store. He therefore brought it down to the workshop again ready for Terry to inspect and hopefully repair the next day.

While upstairs in the morning, Nigel B brought the graduable steam brake down to the workshop for cleaning. While he was stripping it down, he found that the bolt that secures the operating handle had sheared at a point where a small taper pin fits through the casting and also the bolt. The inboard portion of the bolt was removed quite easily with a screwdriver, and then Bill took both pieces of bolt to machinist Barry Nesom who soon made a new one - many thanks again to Barry. While reassembling, Bill had to drill through the bolt via the small hole in the casting so as to fit the small taper pin. The attached photos illustrate the problem and the repair.

 

63395 graduable steam brake showing sheared bolt that secures operating handle and new replacement bolt - Ian Pearson.

63395 graduable steam brake showing new handle securing bolt, and new hole for taper pin - Ian Pearson

63395 graduable steam brake re-assembled with new handle securing bolt - Ian Pearson

Meanwhile, Nigel Hall continued with painting a coat of red gloss on the eccentrics, then black primer on the reverse side of the backplate, the steps and tea tray. He also fitted the left hand top piece of backplate to the boiler and also the mount for the steam pressure gauge. Bill made up a hose connection plate to fit onto the the right hand side clack box so that it could be secured the following day. Ian completed checking and replacing trimmings around the lifting links, cleaned four boiler top doors, and assisted with some of the pipe work and steam brake work.

Wednesday was a dull dismal morning but with the passage of a cold front, there was bright sunshine (but colder) after lunch. Steve Hyman, Ed Bolam, Ian Pearson and Terry Newman were present. Steve and Ed removed the water hose connector from the clack box that was put on yesterday. It required a new gasket and some spacers. Back at the workshop Steve made a new gasket, found 4 spacers and refitted the connection to the clack box. Ed then assisted Terry splitting the vacuum and steam heat pipes to manage the straightening more easily using the oxy gear. By lunchtime, and with some assistance from Ian, both pipes were straight. Ian and Ed then set up a pair of trestles where Ed spent the rest of the day cleaning the pipes ready for painting by Nigel H. There is still one pipe section to clean though. Terry, after obtaining flux and brazing rods from the MPD, repaired the crack in the steam brake pipe. All boiler doors and plugs were taken down to the boiler shop ready for boxing up the boiler and filling it with water, with the intention of a warming fire being lit on Sunday evening prior to a test steaming before the boiler inspector's visit. Ian and Steve spent spent all afternoon cleaning between the frames mainly the motion plates.

Four engines were in steam on Wednesday on test - the 9F, Repton, No 825, and Lucie, which was also tested by pulling the 140 ton 9F up the head shunt. Today, Thursday, saw a return to operation, with Repton on the Christmas diner, which had been steam heated by Lucie. The diner will also be running on4/8/9/10/11/15 to 18/31 December 2020 and 1 to 3 January 2021, departing Grosmont at 12.30. The Christmas Re-imagined service will also be operating from Pickering to Levisham to see Santa on 5/6/12/13 and 19 to 24 December. All require online prebooking and are subject to any change in Covid restrictions.

 

24th and 25 November 2020

Ian Pearson reports a fine breezy Tuesday morning with the odd shower in the afternoon at Grosmont, with Bill Dobson, Nigel Bill, Nigel Hall and Ian himself at work.

On Monday, Sean Bowler fitted the four oiler pads which were put to soak last week, to the trailing and driving wheel sets. Most of Tuesday saw Bill and Nigel Bill, with Ian assisting occasionally, setting up the temporary fire grate under the firebox ready for the out of frames steam test, using scrap items from the area east of Deviation Shed, old pieces of steel and the Q6 fire bars. Nigel Hall continued cleaning and painting the frame with red gloss, while Ian was making and fitting trimmings for lubricating the axle horn cheeks. He also made a start on remaking other motion trimmings.

63395 and 65894 new safety valve fittings. 25 November 2020 - Steve Hyman

In contrast, Wednesday morning was miserable, damp and dismal, but by the afternoon the sun started to shine and the temperature started to drop. Bill Dobson, Ed Bolam, Steve Hyman, Nigel Hall, Terry Newman and Ian were present.

Bill and Ed spent most of the morning completing the building of the temporary fire grate, with a little assistance from Ian and Steve.

Steve's main job was on the J27, servicing the blower valve. After removing the spindle, the head had a very slight groove around it. This was skimmed in the lathe by Barry Nesom. Steve couldn’t clean the seat of the valve as there wasn't a holding tool suitable to clean up the valve seat. So he reassembled the valve and refitted it. Chris Henwood and Barney have been made aware of what has been done up to date.

Ian continued checking and making trimmings on the righthand lifting links, but there are still quite a few to do. Terry repaired three copper pipes which had various defects. He also assessed the state of the bent vacuum and steam heat pipes and how to straighten them. He will have a go next week. Bill collected the refurbished Q6 safety valves for reassembly, having been inspected and passed by Barney (see photo of new parts machined by Barry Nesom - item 3 is for the J27 safety valve).

After Bill had reassembled them, they were fitted to the boiler. The boiler is getting nearer ready for its steam test! Nigel Hall cleaned and primed the reverse side of the backplate cladding, the foot step and tea tray. Down at the MPD, he cleaned and painted the eccentrics on the driving wheel set with red oxide primer.

After lunch round the stove, and also on the bench near the door with Zak the dog, Ed cleaned all remaining oil and dirt from the Q6 spring hangers and horn stays.There was no water later in the afternoon in Deviation shed as it has been isolated due to cold weather precautions.

However, Barney has given an assurance that Deviation Shed can still have its supply while the coaling stage and washing out hose is isolated but that will have to be sorted next week.

 

17th and 18th November 202-0

Ian Pearson reports yet another dull morning on Tuesday, but brighter later and very mild, for Bill Dobson, Nigel Bill, Nigel Hall and Ian.

Terry Newman also came in later in the morning at the request of Bill to inspect some of the copper pipes which were found with defects when removed.

Sean Bowler and Dave Brown spent all day on Monday shimming the horn guides. Sean was quite happy with all except for the left leading front which was was out of square by 60 thou. Sean reported the situation to Bill on Tuesday and explained that the preferred method to rectify this was to machine the horn guide which Sean would do. This machining wasn’t a quick job though, and at the end of Tuesday there was still a few hours machining left for the next day. Bill spent a lot of time sorting through the old sockets to establish how many Whitworth 3/4 drive socket sizes we have. He found quite a few different sizes and is going to purchase from eBay the sizes we don’t have. Of course, most nuts and bolts on locomotives are Whitworth, and we are always having to beg others down at the MPD for certain size sockets. When we get our set completed they will be put on a rack on the workshop wall for easy access. Nigel Bill cleaned the remaining scale from the foundation ring area after making a special tool to withdraw it. Afterwards he cleaned two springs which belong to the intermediate wheel set and then he cleaned 6 underkeeps which Sean and Dave had dismantled on Monday.

 

63395 left leading horn guides being measured on 18 November 2020 - Ian Pearson

63395 horn guide being machined on 18 November 2020 - Ian Pearson

Nigel Hall came in and cleaned and primed the areas on the frames where the NDT inspection had been done, and also some of the frame area where the firebox fits. Ian and Bill assisted Terry finding the defective copper pipes which are stored upstairs. While discussing what action would be required, Ian spotted a cardboard box with 4 new Q6 underkeeps inside. It’s strange that when searching last week they weren't spotted!

Ian spent a lot of time making some new trimmings to fit into the top of the axle boxes to lubricate between the axle box and the horn guides.

Bill was informed by Barry Neesham that he had completed the machining of the safety valves, so these will be ready for assembly shortly.

Wednesday saw a fine mild morning with heavy showers during the afternoon. Bill Dobson, Nigel Bill, Nigel Hall and Ian were present. It was all system go down at the MPD as Sean and Dave had been requested to remove the large lifting jacks from under the Q6, as they were required at Pickering due to the failure of the jacks at the carriage works. So everyone assisted, as these jacks are very heavy and cumbersome not easy to move on a floor surface covered in compressed oil, grease and dirt.

Before the jacks were removed, four more static standing jacks were sought and fitted leaving the Q6 standing on eight static jacks (see photo). The four lifting jacks were moved with their electric cabinet to the entrance of No 4 Road awaiting transport.

63395 on standing jacks at Grosmont MPD on 18 November 2020 - Ian Pearson

Completing this took up to 1pm, and lunch was had around the stove in Deviation shed. Nigel Hall had an earlier lunch, then got started cleaning and painting the inner frames after the removal of the lifting jacks which meant he could get a coating on both side where the firebox fits.

During the morning, Chris Kelly fitted 2 stepped studs onto the backplate which hold the righthand side of the Tea Tray in position. In the machine shop Barry Neesham milled the left leading front horn guide which took some careful machining (see photo). Nigel Bill completed cleaning the under keeps, while Ed painted the Tea Tray. Ian washed the new found oiler pads in warm lubricating oil, as recommended by Armstrong Oilers, and then put them to soak in lubricating oil in the wheely bin in the MPD machine shop.

Some time during the afternoon was spent looking round the site for a suitable steel beam to put under the firebox instead of timber, for when the boiler is given a steam test. However, advice was that a steel plate sat on top of the timber would suffice. Also during the afternoon, Sean and Dave continued fitting, measuring and shimming the left leading front horn guide (see photo).

 

10th and 11th November 2020

Ian reports a mild dull morning, but brightening up just before lunchtime, greeted Bill Dobson, Nigel Bill and Ian himself at Grosmont on Tuesday.

Nigel and Ian continued cleaning brake gear and spring components. Bill spent some time down at the MPD with Nick Simpson, arranging for the two new stepped studs required for the boiler backplate which hold the tea tray in position. While Nigel was cleaning springs, spring hangers and brake tables, Ian continued to clean the remaining six spring hanger bolts in the washer in the repair shed. However, the nut on the the left driving back spring hanger would not come free: it was stuck solid. After several attempts with a large spanner with an extension on its shaft, it still would not move. So the next couple of hours were spent in Deviation Shed trying to remove it with some heat. It was then soon found that the oxygen bottle was empty (the failing equipment curse strikes again!) so the bottle was replaced. This in itself took quite some time as the full ones are in a compound near the boiler shop. After setting up the burning gear with the new oxy bottle it was time for lunch, which was had in Deviation shed with Zack the dog.

After lunch, Nigel did the monthly Defibrillator checks (all OK) and Bill measured up the boiler so he can order the insulation. He also was in conversation with Barry Neesom regarding the safety valves which are almost ready. Ian and Nigel heated the nut on the spring hanger bolt in Deviation Shed and eventually, after about an hour, it finally came off. After it cooled down, it was cleaned, but, as it was still tight on the threads, Bill took it to Barry Neesham who gave a thump on the sides of the nut while screwed on the bolt and this allowed it to turn freely. While Ian cleaned the last two spring hanger bolts, Bill cleaned the paint from inside frames around the horn guides ready for an NDT inspection. Bill did this and Ian took the necessary photos which have been sent to Barney for the Q6 file. No cracks were detected. Sean Bowler had been working on his own shimming the horn guides on Monday and Tuesday, and his next job will be to measure the new distance between the horn guides against each axle box.

Wednesday was a rather dull cool morning but brighter later on, for the rostered Ed Bolam, Nigel Bill, Nigel Hall, Bill Dobson and Ian Pearson. The day was mainly spent cleaning and painting. Last Friday, Nigel Hall gave the boiler barrel another coat of silver paint, and today the fire box area got the same treatment to finish the job. Ed, Ian and Nigel Bill spent all day cleaning the wheel sets and axle box top plates, in the process and removing oil and water from under the top plates along with the old trimmings which will have to be replaced. Bill investigated the state of the underkeep pads. After dismantling the trailing wheel underkeep, the pad was found to be in very poor condition and, after discussing the issue with Barney, it was decided to change all 8 pads. Bill will be putting an order in for these at Armstrong Oilers shortly. Sean Bowler and Dave Brown have completed the shimming of the horn guides, and have fitted new nuts to the horn stays. Now they are all the same size nuts !

Two minutes silence was held at 11am in Deviation Shed and the MPD on this Remembrance Day.

 
  • «
  •  Start 
  •  Prev 
  •  1 
  •  2 
  •  3 
  •  4 
  •  5 
  •  6 
  •  7 
  •  8 
  •  9 
  •  10 
  •  Next 
  •  End 
  • »


Page 1 of 30